40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 30 02:13:23 2017, gautam, Update, General, DRMI locking attempt BS_OLmods.pdf
    Reply  Fri Jun 30 11:50:17 2017, gautam, Update, General, DRMI locking attempt ITMX_glitchy.png
       Reply  Fri Jun 30 22:28:27 2017, gautam, Update, General, DRMI re-locked DRMI_relocked.pngMICH_OL.pdf
          Reply  Sat Jul 1 14:27:00 2017, Koji, Update, General, DRMI re-locked 
Message ID: 13093     Entry time: Fri Jun 30 22:28:27 2017     In reply to: 13090     Reply to this: 13094
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI re-locked 

Summary:

Reverted to old settings, tried to reproduce DRMI lock with settings as close to those used in May this year as possible. Tonight, I was successful in getting a couple of ~10min DRMI 1f locks yes. Now I can go ahead and try and reduce the noise.

I am not attempting a full characterization tonight, but the important changes since the May locks are in the de-whitening boards and coil driver boards. I did not attempt to engage the coil-dewhitening, but the PD whitening works fine.

As a quick check, I tested the hypothesis that the BS OL loop A2L coupling dominates between ~10-50Hz. The attached control signal spectra [Attachment #2] supports this hypothesis. Now to actually change the loop shape.

I've centered Oplevs of all vertex optics, and also the beams on the REFL and AS PDs. The ITMs and BS have been repeatedly aligned since re-installing their respective coil driver electronics, but the SRM alignment needed some adjustment of the bias sliders.

Full characterization to follow. Some things to check:

  • Investigate adn fix the suspect X-arm ASS loop
  • Is there too much power on the AS110 PD post Oct2016 vent? Is the PD saturating?

Lesson learnt: Don't try and change too many things at once!

GV July 5 1130am: Looks like the MICH loop gain wasn't set correctly when I took the attached spectra, seems like the bump around 300Hz was caused by this. On later locks, this feature wasn't present.

Attachment 1: DRMI_relocked.png  64 kB  Uploaded Fri Jun 30 23:29:31 2017  | Hide | Hide all
DRMI_relocked.png
Attachment 2: MICH_OL.pdf  38 kB  Uploaded Fri Jun 30 23:29:58 2017  | Hide | Hide all
MICH_OL.pdf
ELOG V3.1.3-